A researcher wishes to estimate the proportion of adults who have​ high-speed Internet access. What size sample should be obtained if she wishes the estimate to be within 0.02 with 90​% confidence if
​(a) she uses a previous estimate of 0.42?
(b) she does not use any prior estimates?

Answer by Boreal(15235)

interval is e=0.02=1.645*SE, SE=sqrt(.42*.58/n)
square both sides
0.0004=2.706*(.42*58/n)
0.0004n=0.6592
n=1647.95 or 1648
------------------------
use 0.5 and sqrt term is 0.25/n
0.0004=2.706*.25=0.6765
1691.25 or 1692.